Bock Show Me How Calculator Printem Dividing Partnership Income Tyler Hawes and Piper Albright formed a partnership, investin help me plz
Answer #1

Hawes

Albright

a

$147500

$147500

b

$118000

$177000

c

$118000

$177000

d

$145000

$150000

e

$143500

$151500

Part A

In case of no agreement, profit and loss should be distributed equally

295000/2 = 147500

Part B

Ratio of original capital ratio =120000:180000 = 2:3

Hawes = 295000*2/5 = 118000

Albright = 295000*3/5 = 177000

Part C

Hawes

Albright

Interest (5%)

6000

9000

Remaining 280000(295000-6000-9000) in 2:3

112000

168000

Net income

118000

177000

Part D

Hawes

Albright

Salary allowance

40000

45000

Remaining 210000(295000-40000-45000) in 1:1

105000

105000

Net income

145000

150000

Part E

Hawes

Albright

Interest (5%)

6000

9000

Salary allowance

40000

45000

Remaining 195000(295000-6000-9000-40000-45000) in 1:1

97500

97500

Net income

143500

151500